• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 03:01:45 on September28,2005

    Entry number 156308

    keyword=Q2 scan studies part II

    Part II: Q2 scanner with various magnet settings.

    Now we will vary Q3. The sweeper remains "on". The beam is
    still 10 uA.

    Normal Q3 current is L = 1281.7 Amps, R = 1279.8 Amps
    Start the scan with -10% = 1153.5, 1151.8

    Run 5430 -- Q3 at -10%. Aha ! Shoulder changed a lot, and now the right HRS
    has a shoulder too. See fig 1.

    I like this !

    Run 5431 -- Q3 at -5% (L = 1217.6 A, R = 1215.8 A)

    Next, go for +10% to see a dramatic change hopefully.

    Run 5432 -- Q3 at +10% (L = 1409.8 A, R = 1407.8 A)
    Aha again ! Shoulders all gone. See fig 2.

    Run 5433 -- Q3 at +5% (L = 1345.8 A, R = 1343.8 A)
    After this we cycle Q3 and put it back to normal current.

    My happy (and tentative) onclusion is that you can get whatever shoulder you want,
    on either spectrometer, and you can also make
    them disappear, by dialing Q3. It is presumably an optics effect,
    and not "photons, heavy nuclei, backgrounds, or instrumentation
    difficulties" which we had worried about.

    Another conclusion: the scanner device is wonderful.
    One might still be worried that the two HRSs are not identical,
    but I think that is already known and accepted.


    FIGURE 1

    FIGURE 2